OpenStudy (anonymous):

one Solve the equation -1/3+W=5 and 3/8

OpenStudy (anonymous):

really need help

OpenStudy (ranga):

-1/3 + W = 5&3/8 add 1/3 to both sides W = 5&3/8 + 1/3 = 5 + 3/8 + 1/3 = 5 + (9 + 8)/24 = 5 + 17/24 = 5&17/24

OpenStudy (anonymous):

can u help me on another one

OpenStudy (ranga):

I can assist you if you do the work. If you learn how to do one you can do all the rest yourself. But if I just give answers then there won't be any learning and you will have to keepo asking for answers.

OpenStudy (anonymous):

-2 3/7 +b 1/7

OpenStudy (anonymous):

okay

OpenStudy (ranga):

the question is incomplete. there is no equation.

OpenStudy (anonymous):

-2 3/7 +b= 1/7

OpenStudy (anonymous):